# Autoregressive Action Sequence Learning for Robotic Manipulation

### v4

@@ A-B Hyper-parameters and Implementation Details
In this section, we provide a full list of hyper-parameters in Figure A1, Figure A2, and Figure A3 for Push-T, ALOHA, and RLBench, respectively, along with comments on selected hyper-parameters to provide additional implementation details.
@@
Model. The mlp size denotes the hidden feature dimension of the MLP network within the standard multi-head attention operation. The number of latents refers to the number of Gaussians for the Gaussian mixture distribution used to decode continuous actions. The backbone denotes the network used to extract the vision features. We use the ResNet50 for Push-T and ALOHA, and Multi-View Transformer (MVT) for RLBench, identical to the ones used in Diffusion Policy, ACT, and RVT2.
@@
Action Sequence. The horizon refers to the number of actions predicted at each step, while the number of action steps indicates how many actions are actually executed, with the remainder discarded. We adopt the same horizon and action steps as state-of-the-art methods. In Push-T, the chunk size for both high- and low-level actions matches the horizon, meaning all high-level points are predicted in one chunk, followed by all low-level points. Yet, interestingly, as shown in Figure 9, combining these two chunks into a single-step prediction degrades performance. For RLBench, which uses the next key end-effector pose as the control interface, there is no need for high-frequency actions, so neither the horizon nor action steps apply. Instead, low-level robot movements are generated using RLBench’s built-in RRT planner. We use a chunk size of 2 for binary gripper states and a chunk size of 1 for end-effector positions and rotations. For example, ARP first predicts the roll, followed by pitch and yaw of the rotation Euler angle. We follow the strategy of RVT-2 to predict coarse positions and then refine them by zooming into the images (with updated vision features) to obtain more accurate positions. The end-effector positions are predicted in 2-d, and the 3-d positions are derived from the 2-d coordinates of each viewpoint.
Figure A1: Overview of SoTA solutions on Push-T, ALOHA, and RLBench. Diffusion Policy (DP) [13] iteratively subtracts Gaussian noises from noisy actions. The transformer network predicts the Guassian noise at each step. Action Chunking Transformer (ACT) [10] is a VAE architecture that predicts actions directly from images and Gaussian noises. RVT-2 [16] is a hybrid and more complex model, but it is trained directly with behavior cloning and it does not require a generative framework such as diffusion or VAE.
@@
Train & Eval. The observation $2\times 96\times 96\times 3$ represents 2 frames of RGB images, each with a resolution of 96x96 pixels. For RLBench, the observation $4\times 128\times 128\times 4$ refers to RGBD images (with one depth channel) at 128x128 resolution, captured from 4 cameras. In ALOHA, the maximum evaluation steps of 400 and control frequency of 50Hz indicate an evaluation time limit of 8 seconds. LAMB refers to the large batch optimizer. We use the same number of training steps, evaluation frequency, optimizer, learning rate, and learning rate scheduler as used by the SoTA solutions.
@@
A-C Discussion on Action Chunking
@@
We propose two main reasons. First, as has been explored in other imitation learning works, using expert data means that the dataset often lacks information on how to recover from errors, which means that predictions grow worse over time. Using longer action chunks effectively shortens the time horizon. However, we find that action chunking still has noticeable benefits even when the state is well-covered, such as in the Push-T environment. Additionally, this problem becomes less severe as the dataset grows – when the prediction error goes to zero, so does the effect of error recovery.
@@
The second and perhaps stronger explanation is that if the demonstrations are non-Markov, the Markov policy that maximizes single-step accuracy is *not necessarily the optimal policy*. This is true even even if the demonstration policies are optimal, and even in the limit as data and model capacity become infinite. This is because the state occupancy measure is not convex with respect to the policy, so linear combinations of policies can lead to state distributions that are not linear combinations of the demonstration state distributions. This can be address either by learning a non-Markov policy, or by learning a Markov policy that imitates the desired state distribution rather than the demonstrations.
@@
RVT2 66.0 96.0 96.0 88.0 92.0 35.0 80.0 69.0 100.0 99.0
ARP (Ours) 68.0 99.2 94.4 85.6 98.4 35.2 55.2 76.8 90.4 100.0
ARP+ (Ours) 69.6 98.4 86.4 89.6 92.8 46.4 63.2 80.0 97.6 96.0
Figure A5: Performance on RLBench. We report the success rate for each task, and measure the average success rate and rank across all tasks. ARP+ shares the same network definition with ARP but has more layers. The MACs / parameter sizes of RVT-2, ARP, ARP+ are 72.1M/57.1G, 71.9M/56.2G, and 73.8M/57.4G, respectively. ARP performs comparably or outperforms RVT-2 on all tasks. Note that RVT-2 requires current timestep as input, and ARP models do not use timestep.
@@
Figure A6: Demonstrations of all tasks in Push-T, ALOHA, and RLBench. We provide visualizations of key frames from a single episode of Push-T and ALOHA, with the frame order indicated at the bottom right. For RLBench, we visualize one language variant for each task. RLBench features over 100 task variants specified through natural language commands [14], such as "open [pos] drawer" where pos is selected from top, middle, bottom, and "stack [num] [color] blocks", where num ranges from 2, 3, 4, and color is chosen from a palette of 20 colors.
A-D Comparison with RVT-2
Architectural difference between RVT-1 and ARP. Our ARP shares the same visual-language encoder, i.e., Multi-View Transformer (MVT) with RVT-2. Three notable architectural differences between ARP and RVT-2 are:
  1. 1.
    RVT-2 employs a two-stage approach. In the first stage, RVT-2 predicts a coarse end-effector pose from images of the entire workspace. New images are then captured at the predicted coarse pose location, providing finer visual details of the surrounding area. In the second stage, RVT-2 uses these detailed images to predict a fine-grained end-effector pose, which serves as the final output. To implement this, RVT-2 utilizes two MVT encoders—one for coarse inputs and another for fine inputs—along with two separate policy networks for coarse and fine predictions. Each policy network is composed of CNNs and MLPs.
    Similarly, our ARP adopts this two-stage approach for RLBench tasks. We also employ two MVT encoders for coarse and fine inputs, respectively. However, unlike RVT-2, ARP uses a single autoregressive policy network. During the coarse stage, the input visual tokens to this network are from the coarse MVT encoder, while during the fine stage, the input visual tokens come from the fine MVT encoder.
  2. 2.
    RVT-2 uses “Location Conditioned Rotation”, a handcrafted MLP that predicts gripper rotation conditioned on the gripper’s translation. In contrast, ARP achieves a similar effect through autoregression, eliminating the need for a manually designed component.
  3. 3.
    RVT-2 directly upscales the visual features from the MVT encoder to predict pixel coordinates. In contrast, ARP upscales the multiplication of the predicted token (generated through autoregression) with the MVT visual features. Our approach is more aligned with sequence learning, as shown in Figure 7.
We remark that all three architectural differences pertain specifically to the autoregressive policy and are not part of the visual-language encoder.
Why not use timestep as model input. In Figure 8, the RVT-2 result is obtained with the current timestep as input, while our ARP models do not include timestep in their input. We do not use timestep as input for two reasons. First, during training, RVT-2 does not utilize the correct timestep due to an implementation bug in its data loader. This issue can be traced down in the RVT-2 codebase at L239, L265, L392 of rvt/utils/dataset.py. Specifically, the timestep of a frame varies depending on when it is inserted into the replay buffer, which we believe is an unintended behavior by the authors of RVT-2. Accurately emulating this behavior, or migrating our implementation to use this data loader would require substantial engineering effort. Moreover, this data-loader has been noted by other researchers for being confusing, as highlighted in this GitHub discussion (link) and even the code comments from RVT (link). Therefore, we have opted to train ARP without using time-step information.
Second, and more importantly, the timestep in RLBench does not correspond to physical time but rather to the number of macro-steps executed so far in the current episode. Each macro-step consists of three stages: (1) predicting the next end-effector pose and gripper action using a policy (e.g., RVT-2 or ARP), (2) planning a trajectory to reach the predicted pose using RRT, and (3) executing the trajectory and gripper action. As noted by the authors of RVT-2 and PerAct in these GitHub discussions (link1, link2), “time could matter as it informs the network of the current stage of the task”. However, in the real world, the autonomous robot must learn to infer the task stage solely from visual inputs, as there will be no oracle that will be telling the robot the current stage of the task. This capability is critical for scenarios where the robot must interact with the physical world continuously without requiring controlled resets to an initial stage with a timestep of 0, such as placing objects in predefined regions. For these reasons, we have chosen to exclude timestep information when training ARP models.
Dive deeper into timestep and sampling strategy. As we mention above, existing methods for RLBench, such as PerAct, RVT, and ARM are impacted by a flawed training data-loader implementation. The behavior of this data-loader is two-fold:
  1. 1.
    The sampling rate of keyframes gets increased. Note that keyframes are the frames where the gripper stopped, which are provided by RLBench.
  2. 2.
    The timestep of a frame depends on when it is inserted into the replay buffer (randomized).
As a result of its unclear behavior, it is difficult to study the impact of timestep or sampling strategy in training RLBench models. To resolve this issue, we implemented our RLBench dataloader from scratch. Our dataloader has a simple implementation and straightforward behaviors:
  1. 1.
    It only loads keyframes.
  2. 2.
    It provides a correct, non-randomized timestep, if timestep is required as input.
Table A4 compares the results of RVT-2 and ARP on different frame sampling strategies and timestep configurations. The “original sampling + randomized timestep” represents the official implementation of RVT-2. We have some surprising observations:
  1. 1.
    If we remove the randomized timestep, then the performance of RVT-2 drops significantly (81.4 to 77.0). This indicates that the timesteps, due to the unintended randomization, may serve as a regularization instead of providing extra information.
  2. 2.
    The original implementation of RVT-2 requires the timestep as an input. However, we find that by updating the dataloader, we can achieve comparable results without it (81.6 vs. 81.4). This indicates the importance of keyframes in RLBench training, and also verifies the effectiveness of our implementation.
  3. 3.
    If correct timesteps are provided during training, the performance of both RVT-2 and ARP drops drastically (81.4 to 74.1, 81.6 to 77.8). This indicates that the correct timestep is actually harmful to RLBench model trainings, contrary to previous believes. This can be explained by seeing timestep as an information leakage for task stage.
TABLE A4: Impacts of sampling strategy and timestep on RLBench models. The original data-loader of RVT-2 is adopted from prior works, including RVT, PerAct, and ARM. However, this implementation has been noted by researchers as being confusing (discussion, comment). Moreover, as detailed in Appendix A-D, the original data-loader randomizes the timestep of training samples, an unintended behavior by the previous authors. In contrast, we propose a simplified yet equally effective approach that only samples keyframes, which enables the study of the impacts of timestep and provides a solid foundation for future research.
Sampling strategy Timestep (train) Success rate Model
Original Randomized 81.4
None 77 RVT-2
Keyframes Only (Ours) None 81.6
Correct 74.1
None 81.6 ARP
Correct 77.8
Figure A7: Step-by-Step Explanation of Attention Interleaving. We provide a video version of this figure Video/attention-interleaving-tour.mp4 in the supplementary.
It is important to note that the potentially flawed design of this data-loader originates from the early work, C2FARM, rather than RVT-2. This legacy design was subsequently adopted by PerAct, RVT, and RVT-2 to ensure fair comparisons by maintaining consistency in the training data distribution. We hope our released implementation of this simplified data-loader and ARP can be helpful for future research in RLBench or similar robotics environments.

### v5

@@ A.2 Hyper-parameters and Implementation Details
In this section, we provide a full list of hyper-parameters in Table A1, Table A2, and Table A3 for Push-T, ALOHA, and RLBench, respectively, along with comments on selected hyper-parameters to provide additional implementation details.
@@
Model. The mlp size denotes the hidden feature dimension of the MLP network within the standard multi-head attention operation. The number of latents refers to the number of Gaussians for the Gaussian mixture distribution used to decode continuous actions. The backbone denotes the network used to extract the vision features. We use the ResNet50 for Push-T and ALOHA, and Multi-View Transformer (MVT) (goyal2023rvt) for RLBench, identical to the ones used in Diffusion Policy, ACT, and RVT2.
@@
Action Sequence. The horizon refers to the number of actions predicted at each step, while the number of action steps indicates how many actions are actually executed, with the remainder discarded. We adopt the same horizon and action steps as state-of-the-art methods. In Push-T, the chunk size for both high- and low-level actions matches the horizon, meaning all high-level points are predicted in one chunk, followed by all low-level points. Yet, interestingly, as shown in Table 1, combining these two chunks into a single-step prediction degrades performance. For RLBench, which uses the next key end-effector pose as the control interface, there is no need for high-frequency actions, so neither the horizon nor action steps apply. Instead, low-level robot movements are generated using RLBench’s built-in RRT planner. We use a chunk size of 2 for binary gripper states and a chunk size of 1 for end-effector positions and rotations. For example, ARP first predicts the roll, followed by pitch and yaw of the rotation Euler angle. We follow the strategy of RVT-2 to predict coarse positions and then refine them by zooming into the images (with updated vision features) to obtain more accurate positions. The end-effector positions are predicted in 2-d, and the 3-d positions are derived from the 2-d coordinates of each viewpoint.
@@
Train& Eval. The observation $2\times 96\times 96\times 3$ represents 2 frames of RGB images, each with a resolution of 96x96 pixels. For RLBench, the observation $4\times 128\times 128\times 4$ refers to RGBD images (with one depth channel) at 128x128 resolution, captured from 4 cameras. In ALOHA, the maximum evaluation steps of 400 and control frequency of 50Hz indicate an evaluation time limit of 8 seconds. LAMB refers to the large batch optimizer proposed by you2019lamb. We use the same number of training steps, evaluation frequency, optimizer, learning rate, and learning rate scheduler as used by the SoTA solutions.
@@
A.3 Discussion on Action Chunking
@@
We propose two main reasons. First, as has been explored in other imitation learning works, using expert data means that the dataset often lacks information on how to recover from errors, which means that predictions grow worse over time. Using longer action chunks effectively shortens the time horizon. However, we find that action chunking still has noticeable benefits even when the state is well-covered, such as in the Push-T environment. Additionally, this problem becomes less severe as the dataset grows – when the prediction error goes to zero, so does the effect of error recovery ross2011reductionimitationlearningstructured.
The second and perhaps stronger explanation is that if the demonstrations are non-Markov, the Markov policy that maximizes single-step accuracy is *not necessarily the optimal policy*. This is true even even if the demonstration policies are optimal, and even in the limit as data and model capacity become infinite. This is because the state occupancy measure is not convex with respect to the policy, so linear combinations of policies can lead to state distributions that are not linear combinations of the demonstration state distributions (hazan2019provablyefficientmaximumentropy). This can be address either by learning a non-Markov policy, or by learning a Markov policy that imitates the desired state distribution rather than the demonstrations.
@@
Figure A2: Embeddings for Discrete, Continuous, and Pixel-coordinate Actions. Discrete actions are embedded by a simple table lookup on a weight matrix. Continuous actions are embedded with a linear layer. Pixel-coordinate actions are embedded by retrieving the point-wise features at the coordinates on the visual feature maps.
Figure A3: Decoders for Discrete, Continuous, and Pixel-coordinate Actions. For discrete actions, we decode the action embeddings into a categorical distribution with a linear layer followed by a softmax operation. For continuous actions, we decode the embeddings into the parameters of a Gaussian mixture distribution with a linear layer. For the pixel-coordinate actions, we multiply the embedding with a visual feature map or a weight tensor, and convert the result into a 2-d heatmap.
@@
Table A4: Performance on RLBench. We report the success rate for each task, and measure the average success rate and rank across all tasks. ARP+ shares the same network definition with ARP but has more layers. The MACs / parameter sizes of RVT-2, ARP, ARP+ are 72.1M/57.1G, 71.9M/56.2G, and 73.8M/57.4G, respectively. ARP performs comparably or outperforms RVT-2 on all tasks. Notably, ARP+ achieves a 97.3% success rate on the challenging peg insertion task.
@@
RVT2 60.0 100.0 98.7 90.7 88.0 42.7 61.3 60.0 93.3 89.3
ARP (Ours) 69.3 100.0 97.3 85.3 98.7 34.7 52.0 76.0 90.7 100.0
ARP+ (Ours) 74.7 98.7 86.7 89.3 93.3 46.7 62.7 81.3 98.7 93.3
@@
Figure A5: Demonstrations of all tasks in Push-T, ALOHA, and RLBench. We provide visualizations of key frames from a single episode of Push-T and ALOHA, with the frame order indicated at the bottom right. For RLBench, we visualize one language variant for each task. RLBench features over 100 task variants specified through natural language commands (james2020rlbench), such as "open [pos] drawer" where pos is selected from top, middle, bottom, and "stack [num] [color] blocks", where num ranges from 2, 3, 4, and color is chosen from a palette of 20 colors.